OpenStudy (lgbasallote):

integrate: \[\huge \int \frac{(\ln x)^2}{x}dx\]

OpenStudy (ash2326):

\[\int\frac{(\ln x)^2}{x} dx\] Put \{\ln x=t\] We get \[\frac 1 xdx=dt\] So we have \[\int t^2 dt\] I think you can do now, can't you?

OpenStudy (anonymous):

You have to substitute something here..

OpenStudy (anonymous):

You know the derivative of lnx is 1/x so use this advantage..

OpenStudy (lgbasallote):

ohhh of course u-sub \[\huge \frac{(\ln x)^3}{3}?\]

OpenStudy (ash2326):

Yeah, it's the right answer:D

OpenStudy (ash2326):

Don't forget to add the constant

OpenStudy (anonymous):

Oh it is about answer..

OpenStudy (lgbasallote):

oh yeah + C of course haha